### v5

@@ 4.3 National Bargaining Power
In deriving the solution $\lambda_{ji}^{*}$, both countries have a symmetric role and identical prior bargaining power, as per the Nash bargaining axiom. However, countries $i$ and $j$ typically possess asymmetric prior bargaining power. For instance, we could maximize $\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}$ or $\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ given some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013). Without prior knowledge about $\zeta_{ij}$, there is no justification to favor one country over another, resulting in $\lambda^{*}_{ji}$. In a real bargaining context, however, the parameter $\zeta_{ij}$ depends on numerous characteristics between these two, as well as other, countries. Non-economic factors such as military power, population, culture, tradition, legislation, and location, along with alternative optimal criteria, can significantly compromise the symmetry axiom and lead to bargaining solutions that differ substantially from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
We introduce a posterior bargaining power that integrates both the symmetric prior and the data, capitalizing on the results in Theorem 5. To extend the bilateral Nash solution in Theorem 5 to multilateral negotiations, we define
@@
as country $i$’s national bargaining power in the trade system. It is normalized so that all bargaining power sums to one, allowing $\beta_{i}$ to be compared over time. This measure aids multinational negotiations. For instance, when allocating votes in forming or reforming a multinational alliance, using this power can avoid bilateral dialogues and discussions. Despite factors other than production and competitiveness, bilateral trade agreements remain more prevalent than multilateral ones; Yilmazkuday and Yilmazkuday (2014) explored some reasons.
The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ constructs a linear ordering for all countries and maintains a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$. If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$. Specifically, $i$’s bargaining power over $k$ equals the ratio of $i$’s power over $j$ and $j$’s power over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$. Also, we can recover it directly from the relation $\lambda^{*}_{ik}=\beta_{i}/\beta_{k}$.
Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$. Thus, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$. From the production aspect, $\beta_{i}$ declines when companies in country $i$ rely on inputs from partner countries. The more companies depend on partners, the less power the country has to coerce concessions from them. From the consumption aspect, bargaining power increases when the country consumes more imported goods and services, thereby influencing the production of its counterparts. Between the determinants, $g_{i}$ represents the country’s breadth of economy, while $\pi_{i}$ predominantly reflects the depth of the economy, as the matrix $P$ in Eq. (3) already normalizes production volumes. However, the positive correlation between them implies a risk of a domino effect, where a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, leading to further reduction in $g_{i}$. Consequently, $\beta_{i}$ would shrink more than the initial drop in $g_{i}$.
@@ 5 Economic Globalization
Globalization refers to the removal of barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services. A recent example is China’s entry into the WTO in 2001, while the UK’s Brexit in 2020 was a clear rejection of globalization. Conversely, protectionist measures could raise tariffs on imported goods, set import quotas for other countries, and enact stricter government regulations on imports. For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports from other countries (Washington Post, 2018).
Globalization creates winners and losers within a nation but does not automatically compensate the losers, thereby fueling income inequality and political divisions. Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers. This section questions whether the country should retreat from global economic integration without directly attacking any specific country. The objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms. We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
Country $i$’s globalization action triggers adjustments in the entire matrix $P$. Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, with a small change $\Delta P_{ii}$. If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism. As the action has no specific target countries, we assume that whenever $P$ changes the $i$th column, the other columns change proportionally to ensure each row of $P$ still has a unit sum. For example, the change $\Delta P_{ji}$ causes proportional changes $\Delta P_{ji}\sigma_{ji}$ to the $j$th row where:
@@
Consequently, the change $\Delta P_{ii}$ provokes adjustments $\Delta P$ in three levels of fallout. First, the $i$th row in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}$. Secondly, the $i$th column in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}$, due to reprisal reactions, where $\odot$ denotes the element-wise product or Hadamard product and $\Lambda_{i}$ is the $i$th row of $\Lambda$. Lastly, the $i$th column of $\Delta P$ triggers changes in all other columns. We define:
@@
where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with the diagonal vector $\sigma_{ii}\odot\Lambda_{i}$. Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$, and thereby $\Delta P=\Delta P_{ii}M$. So, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M$ as $\Delta P_{ii}\to 0$. We let $M_{i}$ be $M$ with its $i$th column dropped. Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
@@
With the above setting and notations, we have:
@@
The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$. The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimal level because $P_{ii}$ cannot be $1$ or 00. However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may be positive or negative at this optimal level, allowing country $j$ to capitalize on this by making a corresponding negative or positive change in $P_{ij}$, thereby affecting $P_{ii}$. To disincentive other countries from making simultaneous movements, we set $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}$ to be a zero vector. This can be tested by checking if $\pi M_{i}$ is a zero vector, according to Corollary 7.
@@
Given $\Lambda$, Eqs. (17) and (18) for all $i\in\mathcal{N}$ define a dynamical system describing the instantaneous change of $\pi$ with respect to the diagonal of $P$. There are $n$ independent variables in the system since the proportions among off-diagonal elements remain unchanged. Due to the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on the diagonal. For example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% simultaneously; this succession continues until the matrix $P$ converges.
Alternatively, we study the properties of the solution to the first-order conditions. If the solution exists, it satisfies a new counterbalance equilibrium in Eq. (19), as stated in Corollary 8. Eq. (19) places $n-1$ identities to the solution, requiring only one more condition to uniquely identify it. Also, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits some desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, resulting in bilateral authority or trade balance, according to Eq. (6) or Theorem 3, respectively. At the stable solution, any tiny perturbation on the diagonal of $P$ would have a nonsignificant effect on $\pi$, discouraging countries from making small movement. Yet, this solution is optimal for some countries only, and not every country maximizes its competitiveness with this solution. Additionally, the derivatives in Eqs. (17) and (18) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$. Therefore, seeking a far-reaching solution to the first-order conditions might be of limited practical use. Nevertheless, Theorem 6 provides each country with the right direction to steer.
@@
If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
@@
Eq. (19) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
The decision rule on globalization could be as follows: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively. Otherwise, when $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4. In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies. Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and prevent major internal trade disputes. Location and cultural proximity often foster the formation of a PTA or RTA. Besides, Eq. (18) measures the side effects of $i$’s protectionism on the rest of the world, helping select countries for a PTA or RTA. In general, some would benefit from it, whereas others would not.
Finally, Theorem 6 implies a mixed globalization and protectionism strategy, helping $i$ to decide which PTA or RTA to join. Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively. Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq. (17) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$, and the positive and negative contributions are partially offset.
Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as:
@@
In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting. This strategy is better than a pure globalization or protectionism strategy because the right-hand side of Eq. (20) is less than or equal to:
$-\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$, respectively. If most country members of a PTA or RTA are in $\mathcal{N}_{i}^{+}$, then country $i$ should consider participating in the trade agreement. However, country $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply that country $i\in\mathcal{N}_{j}^{+}$, possibly causing additional negotiations if $j$ is already in the PTA or RTA.

### v6

@@ 4.3 National Bargaining Power
In deriving the solution $\lambda_{ji}^{*}$, both countries are assumed to have symmetric roles and identical prior bargaining power, consistent with the Nash bargaining axiom. In practice, however, countries $i$ and $j$ typically exhibit asymmetric prior bargaining power. For example, one could maximize
$\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}\quad\mathrm{or}\quad\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ (16)
for some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013).
In real-world bargaining, however, $\zeta_{ij}$ depends on numerous factors, including non-economic elements such as military strength, population, culture, legal systems, and geographic location. These factors, along with alternative optimal criteria, can significantly violate the symmetry axiom, resulting in bargaining outcomes that deviate from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$. Without prior knowledge of $\zeta_{ij}$, there is no justification to favor one country over another, leading to the symmetric solution $\lambda^{*}_{ji}$, which may provide information about bargaining power.
To address this, we introduce a posterior bargaining power measure that combines the symmetric prior with empirical data, building on Theorem 5. Extending the bilateral Nash solution to multilateral negotiations, we define:
@@
where $\beta_{i}$ represents country $i$’s national bargaining power in the trade system. This measure is normalized so that the sum of all bargaining powers equals one, enabling comparisons over time. It is particularly useful in multilateral context, such as allocating vote rights in forming or reforming international alliances, as it avoids the need for bilateral negotiations. Despite this, bilateral trade agreements remain more common than multilateral ones (e.g., Yilmazkuday and Yilmazkuday, 2014).
The set $\{\beta_{i}\}_{i=1}^{n}$ establishes a linear ordering of countries and preserves the transitivity property of the Nash solution $\lambda_{ij}^{*}$. If $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $\beta_{i}>\beta_{k}$. Specifically, country $i$’s bargaining power relative to $k$ equals:
$\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}=\beta_{i}/\beta_{k},\quad\forall j\in\mathcal{N}.$
Two equally weighted and highly correlated factors determine $\beta_{i}$: production volume $g_{i}$ and competitiveness strength $\pi_{i}$. Consequently, $\beta_{i}$ inherits some of the “controversial” properties of $\pi_{i}$. From the production aspect, $\beta_{i}$ decreases when domestic firms rely heavily on imported inputs, reducing the country’s leverage in negotiations. From the consumption perspective, bargaining power rises when the country imports more goods and services, influencing foreign production. However, the positive correlation between the breadth $g_{i}$ and depth $\pi_{i}$ of the economy introduces a risk of a reverse Matthew effect: a decline in $g_{i}$ can trigger a reduction in $\pi_{i}$, which in turn further depresses $g_{i}$. As a result, $\beta_{i}$ may shrink more than the initial drop in $g_{i}$.
@@ 5 Economic Globalization
Globalization refers to the removal of barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In economic terms, it describes the growing interdependence among countries, driven by trade in goods and services. A notable example is China’s accession to the WTO in 2001, while the UK’s Brexit in 2020 represented a clear rejection of globalization. Conversely, protectionist policies can raise tariffs on imports, impose quotas, and enforce stricter regulations. For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports (Washington Post, 2018b).
Globalization creates both winners and losers within nations, but it does not automatically compensate those who lose, often fueling income inequality and political divisions. Therefore, a more inclusive approach is needed — one that not only maximizes the gains of winners but also mitigates negative impacts on others. The benefits of globalization are also unevenly distributed across countries.
This section explores whether a nation should retreat from global economic integration without directly targeting any specific country. The goal is to enhance national competitiveness rather than simply maximizing the aggregate utility of trade firms. Our analysis focuses on actions that influence the global economy as a whole, while disregarding bilateral trade balances.
5.1 Optimal Level of Globalization
When country $i$ takes an action to adjust its level of globalization, it triggers changes across the entire matrix $P$. The country’s first move affects the diagonal element $P_{ii}$ with a small change $\Delta P_{ii}$. If globalization increases, $P_{ii}$ decreases (i.e., $\Delta P_{ii}<0$); if protectionism rises, $P_{ii}$ increases (i.e., $\Delta P_{ii}>0$).
Since the action does not target specific countries, we assume that whenever column $i$ of $P$ changes, the other columns adjust proportionally to maintain the unit sum of each row. For example, a change $\Delta P_{ji}$ induces proportional changes $\Delta P_{ji}\sigma_{ji}$ within row $j$ where:
@@ 5.1 Optimal Level of Globalization
Consequently, a change $\Delta P_{ii}$ in $P$ produces three levels of adjustments $\Delta P$ in $P$. First, the $i$th row in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}$. Secondly, the $i$th column in $\Delta P$ is $\left[\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}\right]^{\top}$, due to reprisal reactions, where $\odot$ denotes the Hadamard (element-wise) product and $\Lambda_{i}$ is the $i$th row of $\Lambda$. Lastly, the change in column $i$ propagates to all other columns.
We define:
@@
where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with $\sigma_{ii}\odot\Lambda_{i}$ on its diagonal. Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$. Therefore:
$\Delta P=\Delta P_{ii}M,\quad\mathrm{and}\ \mathrm{as}\ \Delta P_{ii}\to 0,\quad\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M.$
Finally, let $M_{i}$ denote $M$ with its $i$th column dropped. Theorem 6 provides the derivative of $\pi$ with respect to $P_{ii}$.
@@
Under the above setting and notations, we have:
@@
The decision rule on globalization could be as follows: country $i$ should pursue further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively. If this derivative is insignificantly different from zero, country $i$ may instead seek trade war targets or trade partnerships as outlined in Sections 3 and 4. In addition, a positive (negative) sign of $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ indicates the country is over-globalized (under-globalized, respectively).
The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$, thereby rebalancing $P$ by $\Lambda$. The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimum level because $P_{ii}$ cannot be $1$ or 0. However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may remain positive or negative at this optimal level, allowing country $j$ to exploit this by adjusting $P_{ij}$, thereby influencing $P_{ii}$. To deter simultaneous strategic moves by other countries, we impose $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$. This condition holds if and only if $\pi M_{i}=\vec{0}_{n-1}$, as stated in Corollary 7.
@@
Given $\Lambda$, Eqs. (17) and (18) for all $i\in\mathcal{N}$ define a dynamical system that describes the instantaneous change of $\pi$ with respect to the diagonal elements of $P$. The system has $n$ independent variables since the proportions among off-diagonal elements remain unchanged. Due to its complexity, numerical algorithms can approximate a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive small adjustments on the diagonal elements. For example, over-globalized countries may increase $P_{ii}$ by 1% at each step, while under-globalized countries reduce $P_{ii}$ by 1% simultaneously. This iterative process continues until both the matrix $P$ and the vector $\pi$ converge. Unlike the convergences discussed in Lau et al. (2022), the diagonal elements do not necessarily converges to the same value.
Alternatively, we examine the properties of the solution to the first-order conditions. If such a solution exists, it satisfies a new counterbalance equilibrium given in Eq. (19), as stated in Corollary 8. Eq. (19) imposes n1n-1 identities to the solution, requiring only one additional condition to uniquely identify it. Moreover, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits certain desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, which correspond to balanced bilateral authority or trade volumes according to Eq. (6) and Theorem 3, respectively. Notably, Eq. (19) reduces to Eq. (3) when $\lambda_{ji}=P_{ij}/P_{ji}$, indicating no globalization or deglobalization is likely occur in the absence of trade imbalance.
@@
If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies:
@@
Eq. (19) holds if all $\lambda_{ij}=\pi_{i}/\pi_{j}$ or all $\lambda_{ij}=P_{ji}/P_{ij}$.
At the stable solution, any small perturbation to the diagonal of $P$ has a insignificant effect on $\pi$, discouraging countries from making minor adjustment. However, this solution is optimal for some countries only and does not necessarily maximize competitiveness for all. Additionally, the derivatives in Eqs. (17) and (18) rely on the exogeneity assumption of $\Lambda$ and proportional rebalancing in $M$. Therefore, pursuing a comprehensive solution to the first-order conditions may have limited practical use. Nevertheless, Theorem 6 provides each country with guidance on the appropriate direction to move.
5.2 Mixed Globalization–Protectionism Strategy
In addition to globalization and trade wars, country $i$ can also form or join preferential trade agreements (PTAs) or regional trade agreements (RTAs) with other economies. Ideally, within a PTA or RTA, the ratios $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should remain balanced to facilitate consensus and prevent significant internal trade disputes. Geographic and cultural proximity often encourage the formation of such agreements. Moreover, Eq. (18) measures the external impact of country $i$’s protectionist policies on the rest of the world, aiding in the selection of suitable partners for PTAs or RTAs. While some countries benefit from these, others may not.
Theorem 6 suggests a mixed globalization–protectionism strategy, guiding country $i$ in deciding which PTAs or RTAs to join. To implement this, we divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on the sign of their components in the vector $\pi M_{i}$:
  • 1.
    $\mathcal{N}_{i}^{+}$: countries with positive components;
  • 2.
    $\mathcal{N}_{i}^{-}$: countries with negative components.
Since $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ is a non-negative vector, we rewrite Eq. (17) as: $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, its contribution to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is negative or positive, respectively, partially offsetting each other. Given $\Delta P_{ii}$, the approximate change in $\pi_{i}$ is:
@@ 5.2 Mixed Globalization–Protectionism Strategy
Under the mixed strategy, country $i$ increases protectionism toward countries in $\mathcal{N}_{i}^{-}$ and deepens globalization with countries in $\mathcal{N}_{i}^{+}$, thereby avoiding the partial offsetting effect. This approach outperforms a purely globalization or purely protectionism strategy because the right-hand side of Eq. (20) is less than or equal to:
$\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, if most members of a PTA or RTA belong to $\mathcal{N}_{i}^{+}$, then country $i$ should consider joining that agreement. However, note that $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply $i\in\mathcal{N}_{j}^{+}$, which may require additional negotiations if country $j$ is already part of the PTA or RTA.